Find a second order homogeneous linear differential equation whose general solution is A tan x + B sin x (A, B constant). [Hint: Use the fact that tan x and sin x are, individually, solutions and solve for the coefficients in standard form.]
Answer #1

Let y(x) = A tanke + B sinx ya dy da A sec²x + 3 cosa and d² y= x2 - A ( 2 secx) (secr toun) - B sina Adding (1) and (3) weY = 4ty + 3 y tann (2 secx+1) - :. 3(25ec’eti) y = 3(y+y) + 2 tann ( 2 sec?se+1) 4 -Y(25ec2x+1) 3y+ 3y + 2 tanx ( 2se